In the equation y=3x-4, a point on the line is what?
Answer by checkley75(3666)

Y=3X-4 TO FIND POINTS ON THIS LINE YOU JUST HAVE TO SELECT ANY VALUE FOR X OR Y AND SOLVE FOR THE OTHER TERM:
SET X=0 & SOLVE FOR Y THUS:
Y=3*0-4
Y=4 OR THE POINT (0,-4) IS ON THIS LINE.
SET Y=2 & SOLVE FOR X THUS:
2=3X-4
3X=2+4
3X=6
X=6/2
X=3 OR THE POINT(3,2) IS ON THIS LINE.